A factory siren has a frequency of 990 Hz.
Calculate the frequency heard by a person
sitting in a car moving at 72 kmh-1 (i) away
from the siren and (ii) towards the siren.
Take the speed of sound in air as 330 ms-1.
1
Expert's answer
2019-05-06T10:11:18-0400

We have

"f_s=990 (Hz); v_o=20 (m\/s); v=330 (m\/s)"

Doppler's effect when source of waves is stationary and observer moving away from source of waves

"f_o=\frac{v-v_o}{v}f_s"

Then

"f_o=\frac{330-20}{330}\cdot 990=930 (Hz)"